This problem says to find all real and imaginary solutions. I don't even know were to start, I know I need to use the quadratic. Thank you
x² + x + square x² + x - 2 = 0

Answer by Alan3354(69443)

x² + x + square x² + x - 2 = 0
--------------
If you mean:
x^2 + x + sqrt(x^2 + x - 2) = 0
x^2 + x = -sqrt(x^2 + x - 2)
Square both sides
x^4 + 2x^3 + x^2 = x^2 + x - 2
x^4 + 2x^3 - x + 2 = 0
This has no real roots. I don't want to spend time on it without being sure it's what the problem is.